OpenStudy (anonymous):

Solve each equation by factoring: 1. 8x^2 + 21= -59x 2. 9m^2 + 48m = -64

OpenStudy (cwrw238):

8x^2 + 59x + 21 = 0 (8x + 3)(x + 7) = 0 can u continue ?

OpenStudy (cwrw238):

9m^2 + 48m + 64 = 0 - this looks like a perfect square as 9 and 64 are perfect squares (3m + 8 )( 3m + 8 ) = 0 m = -8/3

OpenStudy (anonymous):

Yes I can continue, Thank you!!!

OpenStudy (cwrw238):

yw